Grande03:

Hamid published a book of poems. He sold 35 copies in the first week for $10 per copy. He wanted to increase the price of each book to increase his revenue. He found out that for every 50-cent increase, x, in the price per copy, the average number of books he sold decreased by 2 per week. Use this information to create expressions that model the number of books he sold and the revenue he got. The number of books sold is given by the expression ______, and the revenue is given by the expression ______.

Florisalreadytaken:

It's been too long and nobody is replying -- am not really a fan of what am seeing, but i'll try -- so for the 1st one, we're asked to find the number of books sold. the timeframe is undefined, so we'll assume it's \(x \). so we can see that there are 35 copies sold in the 1st week. the number would decrease by 2 per week -- by putting all that in an expression it would be \( 35-x \times 2 \Rightarrow 35-2x\)

Florisalreadytaken:

revenue = number of sold books\( \times \) book price the expressin for the book price would be \( 10+0.50x \) we just put it all together to form a single expression: \[ revenue =( 35-x \times 2 ) ( 10+0.50x) \]

Florisalreadytaken:

we're not asked to simplify that so i think we're done lol.